OpenStudy (anonymous):

find the value of c that makes each trinomial a perfect square. then write the trinomial as a perfect square. the equation ,x^2+11x+c

OpenStudy (anonymous):

Let b be second term then 2(1) b=11 b=11/2 b^2 = 121/4

OpenStudy (anonymous):

\[ x^2 + 11 x + \frac {121}4 = (x+ \frac {11}2 )^2 \]